Find two consecutive odd integers whose product is 23 more than their sum.
Found 2 solutions by checkley77, Alan3354:
Answer by checkley77(12844)

Let x & x+2 be the 2 integers.
x+(x+2)+23=x(x+2)
2x+25=x^2+2x
x^2-25=0
x^2=25
x=sqrt25
x=+-5 ans.
Proof:
5+7+23=5*7
12+23=35
35=35

Answer by Alan3354(69443)

Find two consecutive odd integers whose product is 23 more than their sum.
------------------
x = 1st integer
x+2 = 2nd integer
x+(x+2)+23=x(x+2)
2x+25=x^2+2x
x^2 -25=0
x^2=25
x=sqrt25
x=+-5
--> 5 & 7
--> -5 & -3